The denominator of a fraction is 1 more than the numerator. If the numerator is decreased by one and the denominator is increased by 2, the value of the fraction becomes 2/3. Find the fraction.
Answer by stanbon(48558)

The denominator of a fraction is 1 more than the numerator. If the numerator is decreased by one and the denominator is increased by 2, the value of the fraction becomes 2/3. Find the fraction.
--------------------------
Let the numerator be "x"; then the denominator is "x+1".
-----------
EQUATION:
(x+1)/(x+3) = 2/3
3(x+1) = 2(x+3)
3x+3 = 2x+6
x = 3
x+1 = 4
----------
Ans: 3/4
============
Cheers,
Stan H.